OpenStudy (dr0zier99):

What is the range of the function f(x) = 12 − 3x for the domain {-4, -2, 0, 2, 4}?

OpenStudy (dr0zier99):

{24, 18, 12, 6, 20} {6, 12, 18, 24, 30} {-12, -6, 0, 6, 12} {0, 6, 12, 18, 24}

OpenStudy (dr0zier99):

will medal

zepdrix (zepdrix):

Hey. So you have a collection of `inputs` for this function: {-4, -2, 0, 2, 4} They want to know what `outputs` those will produce when you plug them into the function.

zepdrix (zepdrix):

\[\large\rm f(\color{orangered}{x}) = 12 − 3(\color{orangered}{x})\]Let's plug in our first value from the domain set,\[\large\rm f(\color{orangered}{-4}) = 12 − 3(\color{orangered}{-4})\]So we've plugged in our first input, what is the output? Simplify.

OpenStudy (dr0zier99):

my question says

OpenStudy (dr0zier99):

What is the range of the function f(x) = 12 − 3x for the domain {-4, -2, 0, 2, 4}? {24, 18, 12, 6, 20} {6, 12, 18, 24, 30} {-12, -6, 0, 6, 12} {0, 6, 12, 18, 24}

OpenStudy (dr0zier99):

thats it nothing else

zepdrix (zepdrix):

I know what your question says. You don't need to post it again. I'm asking you a very simple question to lead you to the correct answer. I'm not just going to give it to you.

OpenStudy (dr0zier99):

i got 20 from the formula

zepdrix (zepdrix):

\[\large\rm f(\color{orangered}{-4}) = 12 − 3(\color{orangered}{-4})\]-3 times -4 gives us +12\[\large\rm f(\color{orangered}{-4}) = 12+12\]\[\large\rm f(\color{orangered}{-4}) = 24\] Plugging our first value from the domain into the function, gives us our first range value. { 24, , , , } We don't know the others just yet, but we know that 24 is in that group. It's looking like it's probably A, but you should try another number just to make sure.\[\large\rm f(\color{orangered}{-2}) = 12 − 3(\color{orangered}{-2})\]\[\large\rm f(\color{orangered}{-2}) = ?\]